OpenStudy (anonymous):

rewrite the following in EXPONENTIAL form ln(5a)=x+2

OpenStudy (danjs):

recall e^ln(a) = a

OpenStudy (danjs):

and e^(a+b) = e^a*e^b

OpenStudy (triciaal):

ln = log base e